OpenStudy (anonymous):

Solve. y + 12 = 7 (Points : 4) –5 –4 4 5

OpenStudy (anonymous):

y + 12 = 7 You need to isolate y. So subtract twelve from both sides. y + 12 = 7 - 12 -12 y = ?

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

no, -5 is the correct answer.

OpenStudy (anonymous):



OpenStudy (anonymous):

It's -5. You can plug it in to see for yourself. y + 12 = 7 -5 + 12 = 7 7 = 7 True.

OpenStudy (17cchavez):

y=-5 because you must subtract 12 from both sides and you are left with y=-5